# UniFashion: A Unified Vision-Language Model for Multimodal Fashion Retrieval and Generation

### v1

@@ 5 Conclusion
We have introduced UniFashion, a unified framework that addresses the challenges in multimodal generation and retrieval tasks within the fashion domain. By unifying embedding and generative tasks with a diffusion model and LLM, UniFashion enables controllable and high-fidelity generation, significantly outperforming previous single-task state-of-the-art models across diverse fashion tasks. Our model’s ability to readily adapt to manage complex vision-language tasks demonstrates its potential for enhancing various e-commerce scenarios and fashion-related applications. The findings of this study highlight the importance of exploring the potential learning synergy between multimodal generation and retrieval, and provide a promising direction for future research in the fashion domain.
6 Limitations
This section aims to highlight the limitations of our work and provide further insight into the research in this area. Our model relies on diffusion for multimodal interaction, which means that the composed image generation processes may take longer. In our experiments, we tested the performance of our model on one A100 (80G) GPU. During inference, using 1000 examples from VITON-HD dataset, UniFashion took approximately 3.15 seconds for each image generation. We believe it would be beneficial to explore more efficient sampling methods, such as DPM-Solver++ Lu et al. (2022), to improve the overall efficiency of UniFashion.
@@ Appendix A Basics of Diffusion Models
Given a data point sampled from a real data distribution $x_{0}\in q(x)$, during forward diffusion, $x_{0}$ is gradually “corrupted” at each step $t$ by adding Gaussian noise to the output of step $t$-$1$. It produces a sequence of noisy samples $x_{1},...,x_{T}$. Each step is controlled by:
@@
In the field of image generation, diffusion models operate by progressively denoising a random variable that is sampled from a Gaussian distribution. Latent diffusion models (LDMs) operate in the latent space of a pre-trained autoencoder achieving higher computational efficiency while preserving the generation quality. Stable diffusion model is composed of an autoencoder with an encoder $\mathbb{E}$ and a decoder $\mathbb{D}$, a conditional U-Net denoising model $\bm{\epsilon}_{\theta}$, and a CLIP-based text encoder. With the fixed encoder $\mathbb{E}$, an input image $x$ is first transformed to a lower-dimensional latent space $z_{0}=\mathbb{E}(x)$. The decoder $\mathbb{D}$ performs the opposite operation, decoding $z_{0}$ into the pixel space. When considering a latent variable $z$ and its noisy counterpart $z_{t}$, which is obtained by incrementally adding noises to $z$ over $t$ steps, the latent diffusion models are designed to train the $\bm{\epsilon}_{\theta}(\cdot)$ to predict the added noise $\bm{\epsilon}$ using a standard mean squared error loss:
@@
In the context of our current work, we have a particular focus on the pre-trained multimodal latent diffusion models. For a multimodal conditional generation, given a target image $\mathbf{x}_{0}$, in addition to the textual information, the input condition $\mathbf{y}_{0}$ also contains other constraints such as . The aim is to model the conditional data distribution $q(\mathbf{x}_{0}|\mathbf{y}_{0})$, where $\mathbf{y}_{0}$ contains different modalities prompts. The conditioning mechanism is implemented by first encoding conditional information, then the denoising network $\bm{\epsilon}_{\theta}$ conditions on $y_{0}$ via cross-attention. The label $y_{0}$ in a class-conditional diffusion model $\bm{\epsilon}_{\theta}(x_{t}|y_{0})$ is replaced with a null label $\emptyset$ with a fixed probability during training. At inference time, with a guidance scale $s$, the modified score estimate is further in the direction of $\bm{\epsilon}_{\theta}(x_{t}|y_{0})$ and away from $\bm{\epsilon}_{\theta}(x_{t}|\emptyset)$ as follows:
$\hat{\bm{\epsilon}}_{\theta}(\mathbf{x}_{t}|\mathbf{y}_{0})$ $=\bm{\epsilon}_{\theta}(\mathbf{x}_{t}|\emptyset)+$
$s\cdot(\bm{\epsilon}_{\theta}(\mathbf{x}_{t}|\mathbf{y}_{0})-\bm{\epsilon}_{\theta}(\mathbf{x}_{t}|\emptyset)).$
Appendix B Datasets
We test the effectiveness of UniFashion by experimenting on different tasks including fashion image captioning, cross-modal retrieval, composed image retrieval and composed image generation. Table 6 shows the statistics of the datasets used for two stage in our training process.
For fashion image captioning tasks, we utilize the FashionGen (Zang et al., 2021) dataset. Additionally, to enhance our model’s capability in the CIR task, which involves the ability to retrieve captions for target images, we have annotated images from the training set of Fashion-IQ. Recognizing that manually annotating all the images would be time-consuming and resource-intensive, we draw inspiration from the success of recent MLLM models such as LLaVA in text-annotation tasks, and propose leveraging LLaVA 1.5 (13B) to semi-automatically annotate the dataset. We perform word lemmatization to reduce each word to its root form. Such pre-processing stage is crucial for the Fashion-IQ dataset, as the captions do not describe a single garment but instead express the properties to modify in a given image to match its target. As shown in Fig. 4, by analysis of the captions in Fashion-IQ, we extracted key words that describe clothing information such as color, sleeve, pattern, lace, etc., as prompts for MLLM (LLaVA 1.5). We then instructed the model to generate the corresponding captions referencing words that match the image features, as shown in Fig. 5. After this process, we got the captions for Fashion-IQ dataset. The trained UniFashion from this dataset (Fashion-IQ-cap) can generate captions for images in the evaluation set of Fashion-IQ to assist in the CIR task.
@@ Appendix B Datasets
Appendix C Implementation Details
LLM
During the first phase, due to the flexibility brought by the modular architectural design of BLIP-2, we are able to adapt the model to a broad spectrum of LLMs. In our experiments, in order to effectively utilize the capabilities of the existing MLLM models, we adopted LLaVA-1.5 as the LLM module of the model. Technically, we leverage LoRA to enable a small subset of parameters within UniFashion to be updated concurrently with two layers of adapter during this phase. Specifically, the lora rank is 128 and lora alpha is 256. We utilize the AdamW optimizer with $\beta_{0}$ = 0.9, $\beta_{1}$ = 0.99, and weight decay of 0. The LLMs are trained with a cosine learning rate of 2e-5 and a warmup rate of 0.03. We use a batch size of 32 for the tuned LLMs.
Diffusion Module
Following, StableVITON, we inherit the autoencoder and the denoising U-Net of the Stable Diffusion v1.4. We initialize our denoising U-Net with the weights of the U-Net from the Paint-by-Example and for more refined person texture, we utilized a VAE fine-tuned on the VITONHD dataset from StableVITON. We train the model using an AdamW optimizer with a fixed learning rate of 1e-4 for 360k iterations, employing a batch size of 32. For inference, we employ the pseudo linear multi-step (PLMS) sampler, with the number of sampling steps set to 50.
Appendix D Instruction-Tuning LLMs for Different Caption Style
USER: <Img><queries></Img> + Instruction. Assistant: <answer>.

### v2

@@ 5 Conclusion
We have introduced UniFashion, a unified framework designed to tackle challenges in multimodal generation and retrieval within the fashion domain. By integrating embedding and generative tasks using a diffusion model and LLM, UniFashion enables controllable, high-fidelity generation, significantly outperforming previous single-task state-of-the-art models across various fashion tasks. Our model’s adaptability in handling complex vision-language tasks demonstrates its potential to enhance e-commerce scenarios and fashion-related applications. This study highlights the importance of exploring the learning synergy between multimodal generation and retrieval, offering a promising direction for future research in the fashion domain.
Limitations
In this section, we discuss limitations of our work and offer further insights into research within the fashion domain.
Computational Requirements. UniFashion integrates multiple complex modules, including Q-Former, LLM, and diffusion models, which result in higher computational complexity during training. However, during the inference stage, the computational complexity of UniFashion is comparable to that of current state-of-the-art models. For retrieval tasks, only the Q-Former module is needed to calculate the similarity between the input image or text and the pre-stored candidate features in the database, eliminating the need to utilize the LLM and diffusion model components for inference. For composed image generation tasks, such as fashion design, our model relies on diffusion processes, which may take longer. In our experiments, we tested the performance of our model on an A100 (80G) GPU. During inference, using 1000 examples from the VITON-HD dataset, UniFashion took approximately 3.15 seconds per image generation. We believe exploring more efficient sampling methods, such as DPM-Solver++ Lu et al. (2022), could improve the overall efficiency of UniFashion.
Acknowledgements
We thank the anonymous reviewers for their valuable feedback. This research was partially supported by the grant of HK ITF ITS/359/21FP.
@@ Appendix A Basics of Diffusion Models
Given a data point sampled from a real data distribution $x_{0}\in q(x)$, during forward diffusion, $x_{0}$ is gradually “corrupted” at each step $t$ by adding Gaussian noise to the output of step $t$-$1$. It produces a sequence of noisy samples $\mathbf{x}_{1},\cdots,\mathbf{x}_{T}$. Then, diffusion models learn to reverse the process:
$p(\mathbf{x}_{0:T})$ $=p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t}),$ (14)
$\quad p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})$ $=\mathcal{N}(\mathbf{x}_{t-1};\mu_{t}(\mathbf{x}_{t},t),\sigma_{t}^{2}\mathbf{I}),$
where $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$ is the standard Gaussian distribution and $\mu_{t}(\cdot)$ is the parameterization of the predicted mean. Diffusion models are trained to maximize the marginal likelihood of the data $\mathbb{E}[\log p_{\theta}(\mathbf{x}_{0})]$, and the canonical objective is the variational lower bound of $\log p_{\theta}(\mathbf{x}_{0})$.
@@
Latent diffusion models (LDMs) operate in the latent space of a pre-trained autoencoder achieving higher computational efficiency while preserving the generation quality. Stable diffusion model is composed of an autoencoder with an encoder $\mathbb{E}$ and a decoder $\mathbb{D}$, a conditional U-Net denoising model $\bm{\epsilon}_{\theta}$, and a CLIP-based text encoder. With the fixed encoder $\mathbb{E}$, an input image $x$ is first transformed to a lower-dimensional latent space $z_{0}=\mathbb{E}(x)$. The decoder $\mathbb{D}$ performs the opposite operation, decoding $z_{0}$ into the pixel space. When considering a latent variable $z$ and its noisy counterpart $z_{t}$, which is obtained by incrementally adding noises to $z$ over $t$ steps, the latent diffusion models are designed to train the $\bm{\epsilon}_{\theta}(\cdot)$ to predict the added noise $\bm{\epsilon}$ using a standard mean squared error loss:
@@
In the context of our current work, we have a particular focus on the pre-trained multimodal latent diffusion models. For a multimodal conditional generation, given a target image $\mathbf{x}_{0}$, the input condition $\mathbf{y}_{0}$ could contain different constraints. The aim is to model the conditional data distribution $q(\mathbf{x}_{0}|\mathbf{y}_{0})$, where $\mathbf{y}_{0}$ contains different modalities prompts. The conditioning mechanism is implemented by first encoding conditional information, then the denoising network $\bm{\epsilon}_{\theta}$ conditions on $y_{0}$ via cross-attention. The label $y_{0}$ in a class-conditional diffusion model $\bm{\epsilon}_{\theta}(x_{t}|y_{0})$ is replaced with a null label $\emptyset$ with a fixed probability during training.
Appendix B Implementation Details
LLM
During the first phase, due to the flexibility brought by the modular architectural design of BLIP-2, we are able to adapt the model to a broad spectrum of LLMs. In order to effectively utilize the capabilities of the existing MLLM models, we adopted LLaVA-1.5 as the LLM module of the model. Technically, we leverage LoRA to enable a small subset of parameters within UniFashion to be updated concurrently with two layers of adapter during this phase. Specifically, the lora rank is 128 and lora alpha is 256. We utilize the AdamW optimizer with $\beta_{0}$ = 0.9, $\beta_{1}$ = 0.99, and weight decay of 0. The LLMs are trained with a cosine learning rate of 2e-5 and a warmup rate of 0.03. We use a batch size of 32 for the tuned LLMs.
Diffusion Module
We inherit the autoencoder and the denoising U-Net of the Stable Diffusion v1.4. The weights of the U-Net from Paint-by-Example are used to initialize our denoising U-Net. To achieve more refined person texture, a VAE fine-tuned on the VITONHD dataset from StableVITON is utilized. We train the model using an AdamW optimizer with a fixed learning rate of 1e-4 for 360k iterations, employing a batch size of 32. For inference, we employ the pseudo linear multi-step sampler, with the number of sampling steps set to 50.
@@ Appendix B Implementation Details
Appendix C Datasets
For fashion image captioning tasks, we utilize the FashionGen (Zang et al., 2021) dataset. Additionally, to enhance our model’s capability in the CIR task, which involves the ability to retrieve captions for target images, we have annotated images from the training set of Fashion-IQ. Recognizing that manually annotating all the images would be time-consuming and resource-intensive, we draw inspiration from the success of recent MLLM models such as LLaVA in text-annotation tasks, and propose leveraging LLaVA 1.5 (13B) to semi-automatically annotate the dataset. We perform word lemmatization to reduce each word to its root form. Such pre-processing stage is crucial for the Fashion-IQ dataset, as the captions do not describe a single garment but instead express the properties to modify in a given image to match its target. As shown in Fig. 4, by analysis of the captions in Fashion-IQ, we extracted key words that describe clothing information such as color, sleeve, pattern, lace, etc., as prompts for MLLM (LLaVA 1.5). We then instructed the model to generate the corresponding captions referencing words that match the image features, as shown in Fig. 5. After this process, we got the captions for Fashion-IQ dataset. The trained UniFashion from this dataset (Fashion-IQ-cap) can generate captions for images in the evaluation set of Fashion-IQ to assist in the CIR task. More results can be seen in Fig. 6.
Appendix D Instruction Formats
Due to the disparity in caption styles across different fashion datasets, we employ diverse instructions to fine-tune the LLM, enabling it to generate captions of varying styles. Specifically, the Fashion200K dataset inclines towards providing brief descriptions, the FashionGen dataset is prone to offering professional captions, and in Fashion-IQ-cap, the captions are detailed. Consequently, we have designed distinct instructions for different datasets and tasks, as illustrated in Table 7.
@@ Appendix D Instruction Formats
Figure 6: Caption generation results using our method with images from the Fashion-IQ dataset.
Model Types Task Domain Model Main Structure XMR CIR
Text
Generation
Image
Generation
Cross-modal Retrieval General CLIP (2021) Dual-stream Transfomer
Fashion FashionBERT (2020) Single-stream Transfomer
Multimodal LLM General LLaVA (2023) CLIP, LLM
Composed Image Retrieval General SPRC (2024) CLIP, Qformer
Conditional Diffusion General ControlNet (2023) Stable diffusion
Fashion StableVITON (2023) Stable diffusion
Unified Model General NExT-GPT (2023) ImageBind, LLM, Diffusion
Fashion FAME-ViL (2023) Dual-stream Transfomer
General BLIP2 (2023) CLIP, Qformer, LLM
Unified Model (Ours) Fashion UniFashion CLIP, Qformer, LLM, Diffusion
Table 8: Comparison of different multimodal models. XMR: Cross-modal retrieval tasks; CIR: Compoesd image retrieval task.
Appendix E Visual Results
Figure 3 illustrates the architecture of UniFashion for fine-tuning on the image editing task. Initially, we input the cloth sketch and text guidance into the multimodal encoder. The diffusion model then receives the output from the multimodal encoder, along with the cloth sketches and human features (such as the agnostic mask), to generate the desired images. We compare UniFashion with the MGD (Baldrati et al., 2023b) model for this task. In Fig. 7, we compare the images generated by our approach with the competitor in the VITON-HD (Choi et al., 2021) paired setting. In Fig. 8, we show the generation effects of UniFashion in the VITON-HD unpaired setting. Our method, unlike the MGD method that employs a warping module to generate input sketches, directly uses in-shop garment sketches and is capable of generating images that align more accurately with the provided captions and cloth sketches.
Figure 7: Qualitative comparison on VITON-HD paired test set. From left to right: agnostic-mask image, caption, cloth sketch, MGD-generated image, UniFashion (ours)-generated image and ground truth. Our method is capable of generating images that align more accurately with the given captions and cloth sketch. For optimal viewing, please zoom in.
Figure 8: Qualitative comparison on VITON-HD unpaired test set. From left to right: original image, agnostic-mask image, captions, MGD input sketch, MGD-generated image, UniFashion input sketch and UniFashion (ours)-generated image. Our model is capable of generating images that align more accurately with the provided captions and cloth sketch. For optimal viewing, please zoom in.